OpenStudy (anonymous):

Can someone please help me on this problem? Please? Two machines are packing spaghetti into boxes. The smaller machine can package 400 boxes per hour and the larger machine can package 600 boxes per hour. If the larger machine is on for 2 hours before the smaller machine is turned on, how long after the smaller machine is turned on will a total of 15,000 boxes of spaghetti be boxed? Please, I need help on how to set up this equation) *homework due tomorrow if the Lord will

OpenStudy (anonymous):

If the larger machine is turned on for 2 hours before the small one, 1 200 boxes are already made.

OpenStudy (anonymous):

Next, 1h = 900 boxes

OpenStudy (anonymous):

ahh, 1h = 1000 boxes, sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

So you're missing 13 800 boxes left. 13 800/ 1000 = 13.8 hours.

OpenStudy (anonymous):

thNK YOU

OpenStudy (anonymous):

1,200 + 400t + 600t = 15,000 boxes Can you solve to find the time t?

OpenStudy (anonymous):

YES, THANK YOU

OpenStudy (anonymous):

"Yes" means you can do it ??? If so, do you want me to check your solution?

OpenStudy (anonymous):

YEAH THIS IS WHAT I GOT 1200 + 1000t = 15000 -1200 1000t=13800 t=13.8

OpenStudy (anonymous):

thanks for your help...before you go could you take a look at one more problem?

OpenStudy (anonymous):

t = 13 hr 48 min

OpenStudy (anonymous):

Could you open the new post for each question? That way many helpers can support you.

OpenStudy (anonymous):

okay